OpenStudy (anonymous):

Simplify the expression: (6x^3 − 7x − 4) + (4x^3 + 8x + 3). ^ indicates exponent

OpenStudy (klimenkov):

Use this hint: \(f(x)=ax^3+bx+c, \ g(x)=dx^3+ex+f\) \[f(x)+g(x)=(a+d)x^3+(b+e)x+(c+f)\]

OpenStudy (anonymous):

I have no clue what that is... O_o

OpenStudy (anonymous):

\[6x^3-7x-4+4x^3+8x+3\] \[6x^3+4x^3-7x+8x-4+3\] \[10x^3+x-1\]

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

try combining like terms, like terms are terms that differ on their numerical coefficients.